LBC and MODS checkout 11 Feb 2012 (and flats taken on 12 Feb 2012)

OSA: S. Allanson ISA: O. Kuhn ISS: T. Shih

* Repeated MOS red grating flats with lower VFLAT intensity (VFLAT=8). Flat scripts are:
  • /home/LBTO/dthompson/Kim_Flats/ and are
    • grVflats2.mods34.cal and
    • grVflats2.mods42.cal

* Removed Masks but did not yet load new ones as Tim noticed screws were protruding from some mask frames. He machined the screws and prepared all masks for installing later.

* Observed with MODS while Tim working on this.
  • g191b2b observed in all configurations at airmass 1.1
    • had some difficulty aligning it because the slit was too faint. Increased exposure time to 10 sec. With star nearly centered, enough light illuminated the slit to define it. But when the star was offcenter, the slit could not be centroided. Perhaps using i_sdss would help (but I did not try this).
  • PROBLEM modsAlign requires full path to filename, unlike during last Science run.
  • dual grating, in 90 seconds get ~12000 counts maximum

* Turned ON LBC and took some red data at i, I, z, Y. Way out of focus (because not authorized and LBC collimation model not loaded) but just checked backgrounds. However, moon is just rising.

* Loaded masks while configuring for collecting data for binocular range-balancing. Removed LBC-Red from DX beam so we could try to collect data for the binocular range-balanced collimation model.
  • With John on the phone, we talked about this. MODS AGW is not the one to use. Decided to put it off for later since we are still configured for MODS and still have checkout to do with MODS and LBCs.

* MODS PG0823+546 specphot star.
  • airmass = 1.09
  • PROBLEM During this observation at about 07:20 I noticed laser power was only 0.4mW and setpoint was 0.3mW. How did this change? I checked it earlier (before we changed masks) and it was 1.1mW and setpoint 1mW. Is there something about this script that changed it? Was it because/result of clicking "Lock-On" button on GUI as an alternative to typing imcslock before manually taking another thru-slit image in the acquisition process? Or was it that I stopped and restarted the GUI after the mask exchange?
  • This time centering star in slit did not help modsAlign define the slit. Had to tweak centering by hand/by eye.
  • dual grating, in 300 seconds get about 4000-5000 counts maximum.
  • dual prism, in 180 seconds get about 15000 counts in red and 38000 counts in blue

* Scattered Light in macs0744 field
  • execMODS macs0744_pam30_red.img (in /home/LBTO/SciPrep/20120210Scripts/MODS/ScatteredLight/). This will take a 60-sec exposure at PA=-30 (same as when we saw the scattered light) and through red-only with r-sdss filter. It should be identical to the acq image, but does not need the mask ID.
  • Manually changed laser set point to 1 mW while waiting for collimation and script paused: 09:04 UT
  • Still see scattered light, see mods1r.20120211.0067.fits Pattern may be weaker, but similar shape and position.
    • rotangle = -40 degrees
    • az/el = 295/48 degrees
  • Now running PA=-30 deg script in dual imaging mode. GUICOORDS 07:44:59.321 +39:23:59.240
    • do not see reflection on blue image: mods1b.20120211.0029.fits.
    • increased exposure from 60 sec to 180 sec and ran dual imaging again
      • (note from execMODS -e will also lock IMCS, i.e. it doesn't have to go through instconfig command)
      • see hint of reflection, but weaker than red-only, in red and no reflection in blue. mods1r.20120211.0069.fits and mods1b.20120211.0029.fits
    • back to red-only but with 180 sec
      • (How can we run this without resending the preset? by execMODS -f Instrument to run from Instrument block.)
      • pattern is again extended mods1r.20120211.0070.fits scattered light pattern is different for mirror/dichroic
    • blue-only with 180 sec.
      • execMODS -f Instrument macs0744_pam30_blue3.img
      • no sign of scattered light in resultant image, mods1b.20120211.0030.fits
  • Try PA=150 deg: GUICOORDS 07:44:46.289 +39:30:27.830
    • Red Only: 60-sec: execMODS macs0744_pa150_red.img
    • mods1r.20120211.0071.fits shows no sign of reflection
    • 180-sec #72, still no sign.
  • Try PA=60 deg: GUICOORDS 07:44:39.689 +39:25:26.54
    • execMODS macs0744_pa60_red3.img
    • mods1r.20120211.0073.fits shows no sign of reflection
  • Guide Probe for 150 and 60 deg exposures is close to field, try PA=60 with guide star further off axis
    • execMODS macs0744_pa60_red3gs.img GUICOORDS 07:44:29.796 +39:27:24.98
    • mods1r.20120211.0074.fits shows no sign of reflection
  • Repeat PA=-30 deg red-only to see if same pattern present, az/el = 298/35 rotangle=311.5 deg
    • mods1r.20120211.0075.fits --- reflection still there but pattern slightly shifted.
  • summary: scattered light has something to do with dichroic/mirror (not seen in blue path, only in reflected light) and depends on rotator position.

* Long-Slit acquisiton/test on Q1208+101 which is near moon, 18.4 deg from moon.
  • See how guiding/collim works and practice LS
  • guide star has r_mag=15.7 and b_mag=16.1 in SkyCat catalog "USNO at ESO", actually looks brighter than star for MACS PA=-30 (for that star r_mag=15.5 b_mag=16.0, but elevation was much lower).
  • elev=66.5 --- *there is a reflection in this field too! in upper right quadrant - see mods1r.201200211.0077.fits
  • modsAlign worked well
  • execMODS q1208.obs to take a 10-min spectrum through the 0.6" slit. Seeing is 0.9" on GCS and 1.1" on DIMM.
  • INSTCONFIG TIMEOUT ERROR What part of instconfig caused the problem? We have seen this before and aborting and then resending the script has worked.
  • I could not have chosen a worse PA for this time, HA=00:45 ... parallactic angle 23 deg, PA=-90 deg...
  • ERROR aborting a script: after first image was read out and saved, and just as the next was starting, I cntl-c'd script to abort, but it did not abort:

  • Run dual imaging script on same field to see how this reflection/scattered light looks with the dichroic & on blue as well as red channels.
    • mods1r.20120211.0080.fits (see reflection, this time even with dichroic it is a band and looks sort of like what is seen with mirror, #77, though sharper) and mods1b.20120211.0032.fits (no reflection seen)

11:09 Reconfiguring for LBCs

11:29 Slewing to WT10_330 for pointing/co-pointing and collimation.
  • dofpia, /first, /x2
    • -13371 0 0 0 0 0 0 0.93 5.27 red
    • -4927 -171 26 -473 -775 82 38 1.25 1.99 blue
  • initial IE/CA corrections are -7.6"/-1.88"

11:43 Slew to focus star for UGC field... will run trial OBs (albeit with bright moon) and test each PA for guide/focus stars.
  • red side is taking 4 tries to converge (z4 is oscillating), while blue converged at first try.
  • temps: g=0.88/a=0.06 on SX and g=0.44/a=-0.08 on DX
  • dofpia finally completed, but in 4 iterations!
  • run focus OB for lbcrangebal: 115043/115006, corrections are 2"/-1.6"

trial gi The galaxy does become saturated but not to the point of bleeding even at the longest exposure times, for g, 120sec and 100 sec for i.

  • For PA=0, there are stars in guide and tech chips and these appear bright enough at g and i, although I only noticed one correction per channel sent.

trial Ur Only at U does the galaxy not saturate. It saturates at r but does not bleed. Longest exp are 370 for U and 200 for r.
  • Still for PA=0 there are stars in the guide and tech chips, but the pupils on the focus chip at U look faint.

*back to focus star (forgot that wasn't necessary) and running dofpia, this time without /x2. And while here, lbcrangebal. Very small offsets except for dRX DX which is >~2".

* 12:32 Run long exposures at Urgi 370/370/360/360 to test limits of bleeding and at PA=60, PA=120 and PA=0 again to check guide/focus stars.
  • Urgi_60a.ob
    • Pupils on U may be too faint and it looks like brighter one is off the edge, see lbcbtec.20120211.123350_2.fits
    • on r, pupil may be ok, but also the bright one is on the edge.
    • i 360sec bleeds. g 360 sec does not bleed, r 370 does not bleed U 370 not saturated.
  • Urgi_120a
    • seeing from DIMM ~ 1.4 and from guide chips ~ 1.2-1.8 and one at 2.5
    • pupils for g and i look ok (CHECK U and r)

13:04 slewing to SA107 from /home/LBTO/Calib_OBs/STDS/SA107_UsUBVrg_rRIYzIV.ob
  • 4-pt dither using 6 filters on B and 7 on R
  • We are already in twilight
  • Converged in 2 iterations
  • 13:13 Running all-filter OB. Finished at 13:30

13:30 Slew to flat field position at HA=-30 (RA=15:50) and DEC=27. Field is OK, not many stars. 13:30 is a little late to go to it.
  • 10k counts at g (1.27 sec) and 18k at i (2sec) in test exposure
  • 18k and 30k at g and i, respectively. (Looks like scaling is off or i is variable?).
  • 14k at U-B and 21k in r-S in test exposure taken at 13:35
  • 22k and 30k respectively at U-B and r-S

13:40 Closing

Dome Dark and Tel at Zenith
  • LBC 25 biases started
  • MODS INAF flats started:
    • execMODS grmosflats.56.FIORE1new.cal
    • execMODS grmoslamps.56.FIORE1.cal
    • execMODS gmosFlatNonino0717.cal
    • execMODS grmosARCNonino0717.cal (power bump which caused workstation and blue ccd controller computer to go down).

Power Bump around 07:30 am MST

MODS Powered OFF around 08:00 am MST in preparation for planned power outage on levels 1 and 2

Following night: 20120212

This was an OSURC science night, but we had to shut around 10pm MST for high winds and clouds and this provided an opportunity to continue taking flats for the INAF programs and for the MODS spectra taken on 20120211.

10:30 UT Run modsWarmStart.pro and then modsWake (which may do the same things already done in WarmStart):

mods1r.20120212.N.fits and mods1b.20120212.N.fits

10:31 UT execMODS gmosFlatNonino0717.cal: (ID532613)
  • r N=3-7 and b N=3-7
10:44 UT execMODS gmosARCNonino0717.cal: (ID532613)
  • r N=8-10 and b N=8-10
10:54 UT execMODS grpixflats.cal
  • timeout error on calmode. AGw X and Y positions are orange. Abort and resend script... works
  • r N=11-15 and b N=11-20
11:21 UT execMODS grmosflats.56.FIORE1new.cal: (ID573929)
  • again, timeout error on calmode and AGw X and Y positions orange. Retry and it works immediately
  • r N=16-20 and b N=21-25
11:44 UT execMODS grmoslamps.56.FIORE1.cal: (ID573929)
  • r N=21-23 and b N=26-28
11:56 UT execMODS ~MODSeng/modsScripts/Calib/bias8K.cal
  • r N=24-33 and b N=29-38

* *The mask for Nonino0647 could not yet be loaded because there was not enough room in the cassette. It will have to be put in at the next mask exchange at the start of the INAF run.*

* Now taking flats for reduction of the spectrophotometric and spectroscopic data taken 20120211:

12:14 UT execMODS prbias.cal (copied to ~LBTO/SciPrep/20120210Scripts/MODSCalib)
  • r N=34-43 and b N=39-48
12:28 UT ./prlamps
  • prism flats using template scripts through 0.3" slit.
  • r N=44-47: ok
  • r N=48-51 and b N=49-52
  • b N=53-56
    • Kr 55 (exptime = 5.0) and Xe 56 (exptime = 20) are both very saturated in lines at the red end ~2223 and 2240! but they have many lines at the blue end while 53 (Ar) is not saturated but has few blue lines. These lines are not saturated in the dual channel mode.

* 13:00 Steve reconfiguring for MODS

13:14 UT prflats1
  • prflats1
    • uses 0.3" slit. There is little room for a wider slit for blue, since exp times are already 1 sec and a ND1.5 or UG5 filter is used. But for Red, though the exptime is 1, Vflat level is 5 so that may be decreased and a wider slit used. For today, take the template flats.
    • b N=57-61 (saturated at X=2200) QTH1 1 sec ND1.5
    • b 62 very saturated QTH1 1 sec UG5 ... problems cntl-cing the meta-script...
  • repeat blue with prflats_b5.cal which halves exposure time (now QTH1 0.5sec and ND1.5) and skips UG5 prflats_b require 0.5 sec and ND1.5, not UG5 also
    • b N=68-72. Have 44000 counts peak at X=2200.
  • edit prflats.cal (dual) to set nimgs=1:
    • b N=73 QTH1 ND1.5 1 sec, peak counts 27000 at X=2109
    • b N=74 QTH1 UG5 1 sec, peak counts 60000 at X=2196
    • r N=52 Vflat 5 GG495 1 sec, peak counts 15000 at X=1782
    • run to take 4 more of each. b: 75-82 and r 53-56
  • edit prflats_r.cal (red only) to set nimgs=1
    • r N=57 peak counts 17000 at X=1782
    • N=58-61 same: Red-Only Vflat 5 GG495 exptime 1

* Prism comparison lamps done through 0.3" slit * Prism through slit flats done but through 0.3" slit (is there any other way?). I modified blue-only template to reduce exposure time and remove UG5 block. * grall will do all grating flats:
  • slitless flats (except grpixflats dual mode which were done for INAF)
  • comparison lamps through 0.6" slit: dual, red and blue-only
  • through slit flats (0.6" slit): dual, red and blue-only
  • modsSleep

* grall started at 14:11 UT or 07:11 MST

* grpixflats_r.cal are OK:
  • N=62-66 use Vflat 5 and ND1.5 1.5 sec exposure time.

* grpixflats_b.cal:
  • N=83-86: Blue-Only Vflat 5 clear 5sec saturates at red end, X > 5430 ! aborted 5th exposure (after 86).
  • N=87-88 :Blue-Only Vflat 10 UG5 2sec also saturates at red end X > 6639 ! aborted 3rd exposure (after 88).

* grlamps_b.cal:
  • b 89 Hg[Ar] ND1.5 2sec brightest line at X=6180 has 34000 counts
  • b 90 Xe+Kr clear 15sec one line (brightest) at X=6980 is saturated
  • b 91 Ar clear 30 sec, brightest lines have 25000 counts.

* grlamps_r.cal
  • r 67 Ne ND1.5 1 sec one line at X=4967 saturated
  • r 68 Xe+Kr ND1.5 1
  • r 69 Ar ND1.5 1

* grlamps.cal

  • b 92 Ne+Hg[Ar] ND1.5 2 sec
  • b 93 same as 90 Xe+Kr
  • b 94 same as 91 Ar

* grflats.cal

* grflats_b.cal

* grfltas_r.cal

* looks like grall finished without any errors. Last image taken at 15:50 UT or 8:50 MST so duration of meta-script grall was 1h40m

* remaining to be done are 5" through-slit flats in grating mode: b, r and dual

-- OlgaKuhn - 11 Feb 2012
Topic revision: r7 - 13 Feb 2012, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback